# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v2

@@ 5.1 Fundamental groupoid functor realization
To simplify, we make use of some well established constructions (classifying space/simplicial set of a small category and its geometric realization), we also make use of one very special case of the co-end construction, called “tensor product of space valued functors over the orbit category” in [DavisLueck, Section 1].
@@
  • step 2: we check that, for each subgroup $H$ of $G$, $(X^{(0)})^{H}$ is in bijection with the units of $\Pi(H)$;
@@
  • step 4: in order to check that, for each subgroup $H$ in $G$, we have that $\Pi_{1}(X^{H})_{|(X^{(0)})^{H}}=\Pi(H)$, we construct an intermediate subspace $Z$ of $X$ which allows to facilitate this computation and we do it;
  • step 5: finally, we check that for each morphism $\alpha$ in $\operatorname{Hom}_{\operatorname{Orb}}(H,K)$, the associated map $X^{K}\to X^{H}$ induces the groupoid morphism $\Pi(\alpha)$ under the identification of $\Pi_{1}(X^{H})_{|(X^{0})^{H}}$ with $\Pi(H)$ and the similar one for the subgroup $K$.
Step 1. The 0-skeleton of any $G$-CW-complex $X$ with fundamental groupoid functor $\Pi$ is directly determined by $\Pi$ itself, more precisely by the units of $\Pi(H)$ for the subgroups $H$ of $G$. Considering an orbit $G/H$ as a discrete topological space with $G$-action, the “identity” functor defines a covariant functor $E\colon\operatorname{Orb}(G)\to\operatorname{G-TOP}$ sending $G/H$ to the discrete topological $G$-space $G/H$. Then we define (and are required to do so)
$X^{(0)}:=E\otimes_{\operatorname{Orb}}\Pi.$
Concretely, by definition of $\otimes_{\operatorname{Orb}}$ in [DavisLueck], this is the (discrete) $G$-space obtained as quotient space
$\bigsqcup_{H\in Ob(\operatorname{Orb}(G))}G/H\times\Pi(H)_{0}/\sim$ (5.1)
@@
Step 2. Let us show that this produces the desired 0-skeleton of the space $X$ to construct. For this, we have to compute the fixed point sets say for the subgroup $H$ of $G$. It is straightforward to see that $gK\in G/K$ is fixed by $H$ if and only if $g^{-1}Hg\subset K$. At the same time, from such an element $gK$, we then get a well defined $G$-map
$\begin{split}G/H&\to G/K\\ uH&\mapsto ugK\end{split}$
which sends the coset $1H\in G/H$ to $gK\in G/K$, where $1$ denotes the unit in $G$. Thus, all the points $(gK,x)\in G/K\times\Pi(K)_{0}$ in $X^{(0)}$ which are fixed by $H$ are identified with a point in the single copy $\{1H\}\times\Pi(H)_{0}$ and therefore the $H$-fixed set of $X^{(0)}$ is a quotient of $\Pi(H)_{0}$. We are done once we have shown that no further identifications occur. This follows from the functoriality of $\Pi$: whenever two $H$-fixed points are identified with each other, they are also identified with a single point in $\{1H\}\times\Pi(H)_{0}$.
@@
$\left(\bigsqcup_{H\in Ob(\operatorname{Orb}(G))}G/H\times|\Pi(H)|\sqcup\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times|\Pi(K)|\right)\Big/\sim$ (5.2)
where the equivalence relation is now generated by a multiple mapping cylinder construction, gluing the ends of the spaces associated to morphisms appropriately to the spaces associated to the objects. Concretely, this is done as follows: let us use the notation $(gH,t,x)_{\alpha}$ to denote an element of the summand $G/H\times[0,1]\times|\Pi(K)|$ associated to $\alpha\colon G/H\to G/K\in\operatorname{Hom}_{\operatorname{Orb}(G)}$. Then we declare for all morphisms $\alpha\in Mor(\operatorname{Orb}(G))$, say $\alpha\colon G/H\to G/K$
$\begin{split}(gH,0,x)_{\alpha}&\sim(gH,\Pi(\alpha)(x))\qquad\forall gH\in G/H,\,x\in|\Pi(K)|,\\ (gH,1,x)_{\alpha}&\sim(\alpha(gH),x)\qquad\forall gH\in G/H,\,x\in|\Pi(K)|.\end{split}$ (5.3)
@@
$W_{0}:=\left(\bigsqcup_{H\in\atop Ob(\operatorname{Orb}(G))}G/H\times\Pi(H)_{0}\sqcup\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times\Pi(K)_{0}\right)\Big/\sim$
@@
Now we define the desired $G$-CW-complex $X$ by identifying in $W$ all the points in $W_{0}$ with their image points under the surjective map $p$, i.e. as the push-out
${W_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X}$
Step 4. We next have to analyze $X$ and compute in particular $\Pi_{1}(X;G)$. In order to do that, we define an intermediate $G$-CW-complex $Z$ in the following way. Note that the space $W$ contains canonically as subspace $\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times|\Pi(H)|$. We therefore obtain as a subspace of $X$ the push-out
${W_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${W_{0}\cup\left(\displaystyle\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times|\Pi(H)|\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z}$
where we glue the union of the $G/H\times|\Pi(H)|$ along their 0-skeleta with identifications. Observe that we obtain the same space by simply attaching the spaces $|\Pi(H)|$ to the already constructed 0-skeleton $X^{(0)}$ as the following push-out
${\displaystyle\bigsqcup_{H\in\mathcal{ALL}}G/H\times|\Pi(H)|^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\displaystyle\bigsqcup_{H\in\mathcal{ALL}}G/H\times|\Pi(H)|\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z}$
In particular it is clear that the 0-skeleton of $X$ is the 0-skeleton of $Z$, i.e. precisely $X^{(0)}$.
@@
Using these considerations, let us now compute for a subgroup $H$ of $G$ the fundamental groupoid $\Pi_{1}(X^{H})|_{X^{H}_{(0)}}$ of the $H$-fixed set. We follow the arguments which lead to the identification of
$(\Pi_{1}((X^{(0)})^{H})|_{X^{(0)}})_{0}=(\Pi_{1}(X^{H})|_{(X^{(0)})^{H}})_{0}=\Pi(H)_{0}.$
Indeed, the $H$-fixed set of $Z$ is given as push-out
${(W_{0})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${(W_{0})^{H}\cup\left(\displaystyle\bigsqcup_{K\in\operatorname{Orb}(G);\atop\,\{gK\mid g^{-1}Hg\subset K\}}gK\times|\Pi(K)|\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X^{(0)})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z^{H}}$
@@
Step 5. Now, each such $gK$ gives rise to $\alpha\colon G/H\to G/K;uH\mapsto ugK$ as above which finally gives rise to a $2$-cell in $X^{H}$ which identifies the class of $\{gK\}\times c$ and of $\{1H\}\times\alpha^{*}(c)$ in $\Pi_{1}(Z^{H})$. As before, the other morphisms in $\operatorname{Orb}(G)$ do not add further relations by functoriality of $\alpha$. Next, the $2$-cells in $|\Pi(H)|$ add precisely the required further relations which imply that $\Pi_{1}(X^{H})|_{X^{H}_{(0)}}=\Pi(H)$. Note that the fact that $\alpha^{*}\colon\Pi_{1}(|\Pi(K)|)\to\Pi_{1}(|\Pi(H)|)$ are groupoid homomorphisms implies that no further relations will be created. We obtain the desired result $\Pi_{1}(X^{H})|_{X^{H}_{(0)}}=\Pi(H)$ with a canonical identification.
@@ 5.2 Construction of a universal space
In the following, given a functor $\Pi\colon\operatorname{Orb}(G)\to Groupoid$ (thought of as an abstract fundamental groupoid functor) we construct a universal space $B\Pi$ such that canonically $\Pi_{1}(B\Pi;G)=\Pi$, but all possible higher homotopy groups vanish. It has the universal property that there is a bijection between algebraic maps between fundamental groupoid functors and homotopy classes of maps between $G$-spaces, compare Proposition 5.7.
@@
Then, for each subgroup $H$ of $G$ we first attach cells $H/H\times D^{k}$ of dimension $k$ to $X_{k-1}^{H}$ in order to make $\pi_{k-1}(X^{H})$ trivial. To actually remain in the world of $G$-CW-complexes we induce these attaching constructions up and attach $G/H\times D^{k}$ to $G\cdot X_{k-1}^{H}\subset X_{k-1}$. Note that these $G$-$k$-cells affect also other fixed point sets, but there will be no change of $\pi_{j}$ for j<k1j<k-1, and $\pi_{k-1}$ can only get smaller, but as we make sure that it is trivial, this property will not be affected. After attaching all these $G$-cells we therefore get a new $G$-CW-complex $X_{k}$ containing $X_{k-1}$ such that $\pi_{j}(X_{k}^{H},x)=0$ for all subgroups $H$ of $G$ and for all $2\leq j\leq k-1$ and for all basepoints $x\in X_{k}^{H}$.
@@
It has the following characteristic property:
@@
The $G$-space $B\Pi$ just constructed contains $X$ as a subcomplex and has the same $2$-skeleton as $X$, and satisfies that $\pi_{j}(B\Pi^{H},x)=0$ for all $j\in\mathbb{N}$ and for all subgroups $H$ of $G$ and all $x\in B\Pi^{H}$.
@@
For every $G$-CW-complex $Y$ and for every natural transformation
$\Phi\colon\Pi_{1}(Y;G)\to\Pi$
there exists, unique up to $G$-equivariant homotopy, a $G$-equivariant cellular map
$\varphi\colon Y\to B\Pi$
such that $\varphi_{\#}=\Phi$.
@@
We begin by defining the map $\varphi$ on the $G$-equivariant 0-skeleton of $Y$ by putting
$\varphi_{|Y_{(0)}}:=\Phi(\{e\})_{|Y_{(0)}}.$
Now, we proceed to define $\varphi$ on the $1$-skeleton of $Y$. For a $G$ 11-cell $c$ of the form $G/H\times[0,1]$ (with isotropy $H$) pick the ordinary $1$-cell $c_{H}:=H/H\times D^{1}$ contained in $Y^{H}$. Then $c_{H}$ defines (if we choose an orientation of it) an element $\gamma\in\Pi_{1}(Y^{H})_{|Y^{H}_{(0)}}$. Define now $\phi\colon c_{H}\to(B\Pi^{H})^{(1)}$ such that this map represents the element $\Phi(H)(\gamma)$ in $\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$. Since $H$ acts trivially on the target, this extends uniquely to a $G$-map on the $G$-cell $c$ with values in the $1$-skeleton of $B\Pi$. Of course, we could have picked a different base cell in the $G$-cell $c$. But because $\Phi$ is a natural transformation, the resulting map is independent of this —up to the choice of the representative in the homotopy class of $\Phi(H)(\gamma)$ which had already to be made anyway. This defines the map $\phi$ on the $1$-skeleton of $Y$.
To extend $\phi$ to $Y^{(2)}$, let $c$ be a GG-22-cell of the form $G/H\times D^{2}$. Pick the corresponding ordinary $2$-cell $H/H\times D^{2}$ (with isotropy $H$) which is contained in $Y^{H}$. Its attaching map $\psi\colon S^{1}\to Y^{H}$ is obviously contractible in $Y^{H}$ and hence trivial (i.e. a unit) in the fundamental groupoid $\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$ (we conjugate with a path in $Y^{H}$ to the 0-skeleton to make $\psi$ represent an element of $\Pi_{1}(Y^{H})_{B\Pi^{H}_{(0)}}$, the triviality does not depend on the choice of such a path). Consequently, $\Phi(H)$ being a map of groupoids, also the image under $\Phi(H)$ of this element of the fundamental groupoid is trivial (a unit). By construction of $\phi$ on the $1$-skeleton, this image element is represented by $\phi\circ\psi$ (up to the chosen conjugation with a path to the 0-skeleton), which is hence contractible in $B\Pi^{H}$. Extend $\phi$ over $H/H\times D^{2}$ using this contraction and then extend it $G$-equivariantly to $c=G/H\times D^{2}$.
@@
By the very construction of $\phi$ on the $1$-skeleton, we have $\varphi_{\#}=\Phi$, because the morphism sets of $\Pi(H)=\Pi_{1}(B\Pi^{H})|_{(B\Pi^{(0)})^{H}}$ are generated by the $1$-cells.
@@
It is immediate from the definition that if $\varphi^{\prime}_{\#}=\varphi_{\#}$, then their restrictions to $Y_{(0)}$ are equal. The construction of the desired $G$-homotopy is now done inductively over the skeleta and follows the pattern in the non-equivariant case and for the construction part, making use of the condition that $\phi_{\#}^{\prime}=\phi_{\#}$ for the extension over the $1$-skeleton and of the vanishing of higher homotopy groups for the further extensions.
@@
Figure 1: Torus with action of $\mathbb{Z}_{2}$
In Figure 1 we represent the torus as a square where we identify opposite sides. Because the fixed-point sets are the aspherical spaces $\mathbb{S}^{1}$ for $\mathbb{Z}_{2}$ and $\mathbb{T}^{2}$ for $\{e\}$, this is a classifying space for the functor $\Pi_{1}(T^{2};\mathbb{Z}_{2})$, isomorphically given as
$\Pi\colon\operatorname{Orb}(\mathbb{Z}_{2})\to Groupoids$
which is defined as follows. We have to specify which groupoid it assigns to each of the two possible $\mathbb{Z}_{2}$-orbits, namely to the trivial orbit $\mathbb{Z}_{2}/\mathbb{Z}_{2}$ and the free orbit $\mathbb{Z}_{2}/\{0\}$. Moreover, in the category $\operatorname{Orb}(\mathbb{Z}_{2})$ there are only two non-identity morphisms, the collapse map $\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\mathbb{Z}_{2}$ and the non-identity bijection $\tau\colon\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\{0\}$ and we have to specify the associated morphisms of groupoids.
Concretely, we define $\Pi$ as follows.
@@
Note that indeed this is precisely the fundamental groupoid of the $\mathbb{Z}_{2}$-space $T^{2}$ with the chosen CW-structure, and by the contractibility of $S^{1}$ and $T^{2}$ it is a model for $B\Pi_{1}(T^{2};\mathbb{Z}_{2})$.
@@
$\begin{split}&\mathbb{Z}_{2}/\{0\}\times[0,1]\times T^{2},\qquad\mathbb{Z}_{2}/\{0\}\times[0,1]\times T^{2},\\ &\mathbb{Z}_{2}/\{0\}\times[0,1]\times S^{1}\times\{v_{1},v_{2}\},\qquad[0,1]\times S^{1}\times\{v_{1},v_{2}\}.\end{split}$
@@
Observe that when $G=\{e\}$ is the trivial group and $\Pi(\{e\})$ is a discrete group $\Gamma$ (a groupoid with only one object) then $B\Pi$ is a standard $CW$-complex with a single 0-cell. It follows that the space obtained in this way is an Eilenberg-Mac Lane space $B\Gamma=K(\Gamma,1)$.

### v3

@@ 5.1 Fundamental groupoid functor realization
To simplify, we make use of some well established constructions (classifying space/simplicial set of a small category and its geometric realization), we also make use of one very special case of the co-end construction, called “tensor product of space valued functors over the orbit category” in [1, Section 1].
@@
  • Step 2: we check that, for each subgroup $H$ of $G$, $\bigl(X^{(0)}\bigr)^{H}$ is in bijection with the units of $\Pi(H)$;
@@
  • Step 4: in order to check that, for each subgroup $H$ in $G$, we have that $\smash{\Pi_{1}\bigl(X^{H}\bigr)_{|(X^{(0)})^{H}}}=\Pi(H)$, we construct an intermediate subspace $Z$ of $X$ which allows to facilitate this computation and we do it;
  • Step 5: finally, we check that for each morphism $\alpha$ in $\operatorname{Hom}_{\operatorname{Orb}}(H,K)$, the associated map $X^{K}\to X^{H}$ induces the groupoid morphism $\Pi(\alpha)$ under the identification of $\Pi_{1}\bigl(X^{H}\bigr)_{|(X^{0})^{H}}$ with $\Pi(H)$ and the similar one for the subgroup $K$.
Step 1. The 0-skeleton of any $G$-CW-complex $X$ with fundamental groupoid functor $\Pi$ is directly determined by $\Pi$ itself, more precisely by the units of $\Pi(H)$ for the subgroups $H$ of $G$. Considering an orbit $G/H$ as a discrete topological space with $G$-action, the “identity” functor defines a covariant functor $E\colon\operatorname{Orb}(G)\to\operatorname{G-TOP}$ sending $G/H$ to the discrete topological $G$-space $G/H$. Then we define (and are required to do so) $X^{(0)}:=E\otimes_{\operatorname{Orb}}\Pi$. Concretely, by definition of $\otimes_{\operatorname{Orb}}$ in [1], this is the (discrete) $G$-space obtained as quotient space
$\bigsqcup_{H\in\operatorname{Ob}(\operatorname{Orb}(G))}G/H\times\Pi(H)_{0}/{\sim},$ (5.1)
@@
Step 2. Let us show that this produces the desired 0-skeleton of the space $X$ to construct. For this, we have to compute the fixed point sets say for the subgroup $H$ of $G$. It is straightforward to see that $gK\in G/K$ is fixed by $H$ if and only if $g^{-1}Hg\subset K$. At the same time, from such an element $gK$, we then get a well defined $G$-map $G/H\to G/K$, $uH\mapsto ugK$ which sends the coset $1H\in G/H$ to $gK\in G/K$, where $1$ denotes the unit in $G$. Thus, all the points $(gK,x)\in G/K\times\Pi(K)_{0}$} in $X^{(0)}$ which are fixed by $H$ are identified with a point in the single copy $\{1H\}\times\Pi(H)_{0}$ and therefore the $H$-fixed set of $X^{(0)}$ is a quotient of $\Pi(H)_{0}$. We are done once we have shown that no further identifications occur. This follows from the functoriality of $\Pi$: whenever two $H$-fixed points are identified with each other, they are also identified with a single point in $\{1H\}\times\Pi(H)_{0}$.
@@
$\Bigg(\bigsqcup_{H\in\operatorname{Ob}(\operatorname{Orb}(G))}G/H\times|\Pi(H)|\amalg\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times|\Pi(K)|\Bigg)\Big/{\sim},$
where the equivalence relation is now generated by a multiple mapping cylinder construction, gluing the ends of the spaces associated to morphisms appropriately to the spaces associated to the objects. Concretely, this is done as follows: let us use the notation $(gH,t,x)_{\alpha}$ to denote an element of the summand $G/H\times[0,1]\times|\Pi(K)|$ associated to $\alpha\colon G/H\to G/K\in\operatorname{Hom}_{\operatorname{Orb}(G)}$. Then we declare for all morphisms $\alpha\in\operatorname{Mor}(\operatorname{Orb}(G))$, say $\alpha\colon G/H\to G/K$
$(gH,0,x)_{\alpha}\sim(gH,\Pi(\alpha)(x)),\qquad$ $\forall gH\in G/H,\quad x\in|\Pi(K)|,$
$(gH,1,x)_{\alpha}\sim(\alpha(gH),x),\qquad$ $\forall gH\in G/H,\quad x\in|\Pi(K)|.$ (5.2)
@@
$W_{0}:=\Bigg(\bigsqcup_{H\in\atop\operatorname{Ob}(\operatorname{Orb}(G))}G/H\times\Pi(H)_{0}\amalg\bigsqcup_{\alpha\colon G/H\to G/K\atop\in\operatorname{Hom}_{\operatorname{Orb}(G)}(H,K)}G/H\times[0,1]\times\Pi(K)_{0}\Bigg)\Big/{\sim}$
@@
Now we define the desired $G$-CW-complex $X$ by identifying in $W$ all the points in $W_{0}$ with their image points under the surjective map $p$, i.e., as the pushout
${W_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X.}$
Step 4. We next have to analyze $X$ and compute in particular $\Pi_{1}(X;G)$. In order to do that, we define an intermediate $G$-CW-complex $Z$ in the following way. Note that the space $W$ contains canonically as subspace $\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times|\Pi(H)|$. We therefore obtain as a subspace of $X$ the pushout
${W_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${W_{0}\cup\bigg(\displaystyle\bigsqcup_{H\in\operatorname{Orb}(G)}G/H\times|\Pi(H)|\bigg)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z,}$
where we glue the union of the $G/H\times|\Pi(H)|$ along their 0-skeleta with identifications. Observe that we obtain the same space by simply attaching the spaces $|\Pi(H)|$ to the already constructed 0-skeleton $X^{(0)}$ as the following pushout:
${\displaystyle\bigsqcup_{H\in\mathcal{ALL}}G/H\times|\Pi(H)|^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\displaystyle\bigsqcup_{H\in\mathcal{ALL}}G/H\times|\Pi(H)|\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X^{(0)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z.}$
In particular, it is clear that the 0-skeleton of $X$ is the 0-skeleton of $Z$, i.e., precisely $X^{(0)}$.
@@
Using these considerations, let us now compute for a subgroup $H$ of $G$ the fundamental groupoid $\Pi_{1}\bigl(X^{H}\bigr)|_{X^{H}_{(0)}}$ of the $H$-fixed set. We follow the arguments which lead to the identification of
$\bigl(\Pi_{1}\bigl(\bigl(X^{(0)}\bigr)^{H}\bigr)|_{X^{(0)}}\bigr)_{0}=\bigl(\Pi_{1}\bigl(X^{H}\bigr)|_{(X^{(0)})^{H}}\bigr)_{0}=\Pi(H)_{0}.$
Indeed, the $H$-fixed set of $Z$ is given as pushout
${(W_{0})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${(W_{0})^{H}\cup\Bigg(\displaystyle\bigsqcup_{K\in\operatorname{Orb}(G);\atop\{gK\mid g^{-1}Hg\subset K\}}gK\times|\Pi(K)|\Bigg)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigl(X^{(0)}\bigr)^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z^{H}}$
@@
Step 5. Now, each such $gK$ gives rise to $\alpha\colon G/H\to G/K$; $uH\mapsto ugK$ as above which finally gives rise to a $2$-cell in $X^{H}$ which identifies the class of $\{gK\}\times c$ and of $\{1H\}\times\alpha^{*}(c)$ in $\Pi_{1}\bigl(Z^{H}\bigr)$. As before, the other morphisms in $\operatorname{Orb}(G)$ do not add further relations by functoriality of $\alpha$. Next, the $2$-cells in $|\Pi(H)|$ add precisely the required further relations which imply that $\Pi_{1}\bigl(X^{H}\bigr)|_{X^{H}_{(0)}}=\Pi(H)$. Note that the fact that $\alpha^{*}\colon\Pi_{1}(|\Pi(K)|)\to\Pi_{1}(|\Pi(H)|)$ are groupoid homomorphisms implies that no further relations will be created. We obtain the desired result $\smash{\Pi_{1}\bigl(X^{H}\bigr)|_{X^{H}_{(0)}}=\Pi(H)}$ with a canonical identification.
@@ 5.2 Construction of a universal space
In the following, given a functor $\Pi\colon\operatorname{Orb}(G)\to\operatorname{groupoid}$ (thought of as an abstract fundamental groupoid functor) we construct a universal space $B\Pi$ such that canonically $\Pi_{1}(B\Pi;G)=\Pi$, but all possible higher homotopy groups vanish. It has the universal property that there is a bijection between algebraic maps between fundamental groupoid functors and homotopy classes of maps between $G$-spaces, compare Proposition 5.7.
@@
Then, for each subgroup $H$ of $G$ we first attach cells $H/H\times D^{k}$ of dimension $k$ to $X_{k-1}^{H}$ in order to make $\pi_{k-1}\bigl(X^{H}\bigr)$ trivial. To actually remain in the world of $G$-CW-complexes we induce these attaching constructions up and attach $G/H\times D^{k}$ to $G\cdot X_{k-1}^{H}\subset X_{k-1}$. Note that these $G$-$k$-cells affect also other fixed point sets, but there will be no change of $\pi_{j}$ for j<k1{j<k-1}, and $\pi_{k-1}$ can only get smaller, but as we make sure that it is trivial, this property will not be affected. After attaching all these $G$-cells we therefore get a new $G$-CW-complex $X_{k}$ containing $X_{k-1}$ such that $\pi_{j}\bigl(X_{k}^{H},x\bigr)=0$} for all subgroups $H$ of $G$ and for all $2\leq j\leq k-1$ and for all basepoints $x\in X_{k}^{H}$}.
@@
It has the following characteristic property.
@@
The $G$-space $B\Pi$ just constructed contains $X$ as a subcomplex and has the same $2$-skeleton as $X$, and satisfies that $\pi_{j}\bigl(B\Pi^{H},x\bigr)=0$ for all $j\in\mathbb{N}$ and for all subgroups $H$ of $G$ and all $x\in B\Pi^{H}$.
@@
For every $G$-CW-complex $Y$ and for every natural transformation $\Phi\colon\Pi_{1}(Y;\allowbreak G)\to\Pi$, there exists, unique up to $G$-equivariant homotopy, a $G$-equivariant cellular map $\varphi\colon Y\to B\Pi$ such that $\varphi_{\#}=\Phi$.
@@
We begin by defining the map $\varphi$ on the $G$-equivariant 0-skeleton of $Y$ by putting $\varphi_{|Y_{(0)}}:=\smash{\Phi(\{e\})_{|Y_{(0)}}}$. Now, we proceed to define $\varphi$ on the $1$-skeleton of $Y$. For a $G$ 11-cell $c$ of the form $G/H\times[0,1]$ (with isotropy $H$) pick the ordinary $1$-cell $c_{H}:=H/H\times D^{1}$ contained in $Y^{H}$. Then $c_{H}$ defines (if we choose an orientation of it) an element $\gamma\in\Pi_{1}\bigl(Y^{H}\bigr)_{|Y^{H}_{(0)}}$. Define now $\phi\colon c_{H}\to\bigl(B\Pi^{H}\bigr)^{(1)}$ such that this map represents the element $\Phi(H)(\gamma)$ in $\Pi_{1}\bigl(B\Pi^{H}\bigr)|_{(B\Pi^{(0)})^{H}}$. Since $H$ acts trivially on the target, this extends uniquely to a $G$-map on the $G$-cell $c$ with values in the $1$-skeleton of $B\Pi$. Of course, we could have picked a different base cell in the $G$-cell $c$. But because $\Phi$ is a natural transformation, the resulting map is independent of this —up to the choice of the representative in the homotopy class of $\Phi(H)(\gamma)$ which had already to be made anyway. This defines the map $\phi$ on the $1$-skeleton of $Y$.
To extend $\phi$ to $Y^{(2)}$, let $c$ be a GG-22-cell of the form $G/H\times D^{2}$. Pick the corresponding ordinary $2$-cell $H/H\times D^{2}$ (with isotropy $H$) which is contained in $Y^{H}$. Its attaching map $\psi\colon S^{1}\to Y^{H}$} is obviously contractible in $Y^{H}$ and hence trivial (i.e., a unit) in the fundamental groupoid $\Pi_{1}\bigl(B\Pi^{H}\bigr)|_{(B\Pi^{(0)})^{H}}$ (we conjugate with a path in $Y^{H}$ to the 0-skeleton to make $\psi$ represent an element of $\Pi_{1}\bigl(Y^{H}\bigr)_{B\Pi^{H}_{(0)}}$, the triviality does not depend on the choice of such a path). Consequently, $\Phi(H)$ being a map of groupoids, also the image under $\Phi(H)$ of this element of the fundamental groupoid is trivial (a unit). By construction of $\phi$ on the $1$-skeleton, this image element is represented by $\phi\circ\psi$ (up to the chosen conjugation with a path to the 0-skeleton), which is hence contractible in $B\Pi^{H}$. Extend $\phi$ over $H/H\times D^{2}$ using this contraction and then extend it $G$-equivariantly to $c=G/H\times D^{2}$.
@@
By the very construction of $\phi$ on the $1$-skeleton, we have $\varphi_{\#}=\Phi$, because the morphism sets of $\Pi(H)=\Pi_{1}\bigl(B\Pi^{H}\bigr)|_{(B\Pi^{(0)})^{H}}$ are generated by the $1$-cells.
@@
It is immediate from the definition that if $\varphi^{\prime}_{\#}=\varphi_{\#}$, then their restrictions to $Y_{(0)}$ are equal. The construction of the desired $G$-homotopy is now done inductively over the skeleta and follows the pattern in the non-equivariant case and for the construction part, making use of the condition that $\phi_{\#}^{\prime}=\phi_{\#}$ for the extension over the $1$-skeleton and of the vanishing of higher homotopy groups for the further extensions. ∎
@@
Figure 1: Torus with action of $\mathbb{Z}_{2}$.
In Figure 1, we represent the torus as a square where we identify opposite sides. Because the fixed-point sets are the aspherical spaces $\mathbb{S}^{1}$ for $\mathbb{Z}_{2}$ and $\mathbb{T}^{2}$ for $\{e\}$, this is a classifying space for the functor $\Pi_{1}\bigl(T^{2};\mathbb{Z}_{2}\bigr)$, isomorphically given as $\Pi\colon\operatorname{Orb}(\mathbb{Z}_{2})\to\operatorname{Groupoids}$ which is defined as follows. We have to specify which groupoid it assigns to each of the two possible $\mathbb{Z}_{2}$-orbits, namely to the trivial orbit $\mathbb{Z}_{2}/\mathbb{Z}_{2}$ and the free orbit $\mathbb{Z}_{2}/\{0\}$. Moreover, in the category $\operatorname{Orb}(\mathbb{Z}_{2})$ there are only two non-identity morphisms, the collapse map $\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\mathbb{Z}_{2}$ and the non-identity bijection $\tau\colon\mathbb{Z}_{2}/\{0\}\to\mathbb{Z}_{2}/\{0\}$ and we have to specify the associated morphisms of groupoids.
Concretely, we define $\Pi$ as follows:
@@
Note that indeed this is precisely the fundamental groupoid of the $\mathbb{Z}_{2}$-space $T^{2}$ with the chosen CW-structure, and by the contractibility of $S^{1}$ and $T^{2}$ it is a model for $B\Pi_{1}\bigl(T^{2};\mathbb{Z}_{2}\bigr)$.
@@
$\mathbb{Z}_{2}/\{0\}\times[0,1]\times T^{2},\qquad\mathbb{Z}_{2}/\{0\}\times[0,1]\times T^{2},\qquad\mathbb{Z}_{2}/\{0\}\times[0,1]\times S^{1}\times\{v_{1},v_{2}\},$
$[0,1]\times S^{1}\times\{v_{1},v_{2}\}.$
@@
Observe that when $G=\{e\}$ is the trivial group and $\Pi(\{e\})$ is a discrete group $\Gamma$ (a groupoid with only one object) then $B\Pi$ is a standard $CW$-complex with a single 0-cell. It follows that the space obtained in this way is an Eilenberg–Mac Lane space $B\Gamma=K(\Gamma,1)$.